# Network-Level Prompt and Trait Leakage in Local Research Agents

### v2

@@ 7.1 Prompt Recovery
ICL vs. Fine-Tuning Across Models. To motivate our choice of ICL over fine-tuning as the primary task-specific approach, we compare their performance across closed-source models (GPT-4o and GPT-4.1 nano) and open-source models (Llama 3.1 8B, Gemma 3 4B, and Qwen 2.5 32B) (Table 2). We create our final training set by combining the training samples from the prompt recovery datasets discussed in Section 6.1. For closed-source models, fine-tuning is performed through OpenAI’s fine-tuning API [54] for 3 epochs with batch size 1. For open-source models, we apply LoRA [29] with rank 16 and . Results show that ICL generally outperforms fine-tuning across nearly all metrics, though models like Llama occasionally favor fine-tuning for functional and domain equivalence. We attribute this to limited sample diversity, which likely causes overfitting in fine-tuned models.
@@ 7.2 Trait Inference
Figure 5 compares similarity scores for selected and unselected traits. Selected traits typically exceed 0.7, with some personas (012, 024, 037, 045) surpassing 0.8, indicating consistent high-risk leakage [10]. Unselected traits generally fall below but often remain above 0.5, showing moderate exposure even when traits are not explicitly embedded. In a few cases (007, 025, 041), unselected traits approach or exceed selected ones, confirming leakage extends beyond explicitly mentioned attributes.
@@ 8 Defense and Realistic Deployment
If a task requires visiting a uniquely identifying domain, the defense should hide that interaction via VPNs or randomized background queries. By contrast, if the information can be obtained without distinctive traces, blocking is preferable; for example, by querying an LLM directly or gathering information from Wikipedia. We thus identify two conceptual strategies for mitigating metadata-based inference attacks: (1) hiding traces, where sensitive activity must occur but is camouflaged with plausible noise; and (2) blocking traces, where sensitive activity is avoided by relying on alternative, less revealing sources.
@@ 8.1 Hiding via Decoy Prompts
Inspired by TrackMeNot [48], for each real prompt the system generates and executes additional decoy prompts that remain within the same topical space but differ in framing, intent, or implied traits. Because OBELS evaluates whether reconstructions preserve intent, domain type, and entity, breaking this alignment is critical; otherwise, an adversary can still recover the user’s functional behavior despite wording changes. For instance, a prompt about “recognizing depression and seeking medical help for a friend” can be paired with a decoy like “compare the effectiveness of yoga versus meditation for improving sleep quality”, which shifts both domain and entity, lowering OBELS score and obscuring the true goal. This ambiguity in observed traffic undermines an adversary’s ability to infer the user’s true objective, attributes, or target entity. We explore a lightweight approach using LLM-based decoy generation and a trait-aware variant that leverages virtual personas to introduce long-term behavioral divergence.
@@ 8.2 Blocking via Alternative Sources
A complementary defense is to prevent sensitive traces from arising in the first place. When a user’s objective can be met without contacting uniquely identifying domains, the agent can instead draw from large multipurpose repositories (e.g., Wikipedia, StackExchange, Reddit) or from the LLM’s internal knowledge. By retrieving information from resources that serve diverse intents or by leveraging the model’s internal knowledge, the resulting traffic no longer maps cleanly to a specific user goal. This “blocking” approach avoids generating distinctive traces altogether, reducing the adversary’s ability to associate domain visits with private attributes.
@@
Table 10 and Table 11 together show that visibility of browsing traces strongly shapes leakage but has little effect on utility. Utility is how much the WRA’s output helps a human reader reach their research goals. We measure it with an LLM using the prompt template in Figure 17. URL-level traces leak substantially more than domain-only traces, making prompt inference easier and more accurate. Reducing visibility (e.g., observing only 80% or 60% of domains) lowers leakage but still exposes useful information. Adding timing metadata to domains yields no consistent gains, suggesting that domain identity is more informative than temporal order. In contrast, the utility of the generated reports remains relatively stable across visibility levels down to 10%. However, beyond this threshold, utility begins to drop, revealing diminishing returns from additional domain coverage. This suggests that many extra domains contribute little to report quality while disproportionately increasing exposure to adversaries. Thus, WRA behaviors that aggressively visit many domains are especially problematic: they create substantial leakage risks while providing only marginal gains in utility.
@@ 9 Discussion and Conclusion
Privacy and Security Implications. Our results show that metadata leakage is an inherent risk for locally deployed web and research agents. We find that design choices—including domain exploration breadth, backbone model capacity, and orchestration logic—directly affect the amount of information exposed. Substituting locally deployable models does not eliminate leakage: domain-level traffic is structurally unavoidable. This shifts the attack vector from content to behavior. Whereas prompt injection and data exfiltration could be mitigated through isolation [77], network-level traces cannot be trivially suppressed. At scale, such leakage enables profiling by advertisers, ISPs, or state actors, raising unique questions about how to balance agent utility with user privacy.
Defense Effectiveness and Limitations. Our study has several limitations. We rely on proxy datasets and synthetic personas with limited session counts; longer-term aggregation or multimodal attacks may amplify risks. Defense evaluation is confined to decoys and blocking, and cost/latency trade-offs are only partially measured. Nonetheless, our results show that decoy prompts and trait-conflicting personas reduce both prompt recovery and trait inference accuracy, while blocking strategies help when tasks can rely on broad, non-unique sources (e.g., Wikipedia) without sacrificing utility. Yet defenses remain partial: adversaries still recover useful signals, and decoys introduce overhead in latency, bandwidth, and behavioral realism. As with prior obfuscation systems, mitigation is meaningful but incomplete; full network-level protections (e.g., VPNs or anonymity systems for agents [59]) remain the only robust safeguard.
Future Directions and Open Challenges. Future work should explore hybrid defenses that combine obfuscation, blocking, and timing perturbations; study long-term aggregation attacks across extended user histories; and integrate formal privacy guarantees such as differential privacy or traffic-analysis resistance directly into agent design. Multi-modal agents introduce additional risks, for example, as passed multi-modal content could reveal even more information [66]. Overall, progress requires viewing AI agents and the web as a coupled ecosystem rather than orthogonal components.
@@ Acknowledgments
The work was partially supported by Schmidt Sciences SAFE-AI program and by NSF 2333965.
@@ Appendix A Ethical Considerations
Our research investigates privacy risks in Web and Research Agents (WRAs) by studying metadata-based prompt and trait leakage attacks. The goal is to better understand how browsing traces may expose sensitive information and to encourage the development of more secure and privacy-preserving agents. By identifying concrete risks and suggesting defenses, we aim to make a constructive contribution to AI safety.
The main stakeholders are end-users of AI agents, who risk exposure of personal traits or prompts; developers and providers, who may need to adapt their systems to mitigate these risks; and the broader public, which benefits from safer deployments of agent technologies. Our experiments used only public datasets (FEDWEB13, SESSION14, DD16) and synthetic personas, ensuring that no real user data or personally identifiable information was involved. Also, no live systems were disrupted.
We recognize that the methods presented could, in principle, be misused by malicious actors. To balance reproducibility with responsibility, we release the full code to support transparency and community evaluation, but note that it operates only on public datasets and synthetic personas. This minimizes direct risk of harm while enabling other researchers to validate our findings and develop stronger defenses.
We believe that the benefits of publishing this research, raising awareness of risks, guiding the design of defenses, and informing regulatory discussions, significantly outweigh the potential harms. By conducting the research within strict ethical and legal boundaries, and by emphasizing mitigations, we ensure that the work contributes constructively to the responsible advancement of AI systems.
@@ Appendix B Open Science
Implementations, evaluation code, prompts, and the synthetic datasets derived from our study will be made publicly available. Code available at: https://github.com/umass-aisec/wra.
@@ C.1 Dataset Details
Prompt Recovery. For all TREC datasets used in the prompt recovery task, we create -DR variants by inputting each original prompt, along with the suggested_rewriting_prompt from [32], to GPT-4.1. These rewritten prompts are used for the Deep Research API runs to match its prompt handling, and all evaluations are scored against the original, unmodified prompts.
TREC FedWeb 2013 Topics111https://trec.nist.gov/data/federated/2013/fedweb13_50topics.xml (FEDWEB13) contains 50 topics originally designed to evaluate federated web search. We concatenate the <description> and <narrative> fields to form a single prompt, and apply GPT-4.1 rewriting to richer instruction-complete variants (FEDWEB13-DR).
TREC Session 2014 Topics222https://trec.nist.gov/data/session2014.html (SESSION14) contains 60 realistic user topic descriptions spanning domains such as travel, health, and education. We use the original description fields as user prompts and their rewritten counterparts (SESSION14-DR).
TREC Dynamic Domain 2016 Truth Data333https://trec.nist.gov/data/domain2016.html (DD16) provides 53 topics focused on interactive information retrieval in evolving domains such as medicine and cybersecurity. Following the same procedure as FEDWEB13, we concatenate the <description> and <narrative> fields and rewrite them with GPT-4.1 (DD16-DR).
@@
Trait Inference Personas. We use SynthLabsAI/PERSONA_subset to simulate benign users who use a web agent over time. The dataset includes 997 synthetic user profiles annotated with 32 traits, grouped into four categories:
Demographic: age, sex, race, ancestry, religion, place of birth, citizenship, income, education, marital status, household type, household language, veteran status, disability, family presence and age, and health insurance
Occupational: employment status, industry, occupation, class of worker, and detailed job description
Psychographic: ideology, political views, and Big Five personality traits
Behavioral: lifestyle, personal time, mannerisms, and defining quirks
For trait inference, we adopt a type-aware scoring strategy tailored to each trait’s structure:
@@
Ordinal traits (e.g., big five scores) are mapped to integer levels and scored by normalized distance on a 5-point scale (e.g., low to high). Closer ratings yield higher similarity scores.
Categorical traits (e.g., gender, marital status) use exact match for single-token values; multi-word categories are scored using the SBERT score to account for semantic equivalence.
@@
In our experiments, we sample 50 personas. For each, including those reserved for ICL examples, we randomly select 5 traits and use GPT-4o (temperature = 0.7) to generate 21–35 trait-revealing prompts that explicitly or implicitly embed selected traits. The produced traces serve as the input for trait inference.
@@ C.2 Agent and LLM Backbone Choice
When browsing is executed locally, the network-level outbound traffic—accessed domains, timing, and payload length—is visible to the adversary; when browsing is server-side (e.g., OpenAI’s Operator), only connections to the provider’s infrastructure are exposed. To study the leakage risks in the local setting, we collect traces using three open-source WRAs, GPT Researcher [21], AutoGen [84], and Browser-Use [43], and one proprietary system, the OpenAI Deep Research API.
GPT Researcher [21]. GPT Researcher is a research agent designed for deep multi-hop investigation. Unlike lightweight browsing agents, it decomposes prompts into sub-questions, iteratively gathers evidence, and synthesizes reports. We run it in report_type=‘‘deep’’ mode, which enables multi-turn exploration over multiple sources. The agent logs visited URLs with timestamps, from which we extract domains. While it does not expose packet-level traces, these logs still capture the trajectory of its structured reasoning.
@@
A distinctive feature of GPT Researcher is its strategic use of different LLMs for different sub-tasks: FAST_LLM (e.g., GPT-4o-mini, DeepSeek-V3-0324) for lightweight operations such as summarization, SMART_LLM (e.g., GPT-4.1, Mistral-7B-Instruct v0.3) for reasoning and report generation, and STRATEGIC_LLM LLMs (e.g., o4-mini, Mistral-Nemo-12B) for higher-level planning such as research planning. Its default configuration assigns OpenAI models, but we also evaluate alternative combinations using local models to assess privacy leakage when deployed outside the proprietary ecosystem.
Browser-Use [43]. Browser-Use is an open-source web agent that automates browsing through a local browser, in contrast to GPT Researcher. By default, it queries Google via search_web, but this frequently triggers reCAPTCHA and stalls sessions. To ensure stability, we configure it to start from Bing and automatically fall back to Bing when blocked.
@@
We further modify its system prompt to require at least five distinct page visits before summarization, preventing it from shortcutting tasks with prior knowledge or superficial searches. After each run, we extract domains and timestamps from the browser’s recorded history to build ordered traces. This setup makes Browser-Use suitable for evaluating how web agents’ lightweight browsing patterns leak information.
AutoGen [84]. AutoGen is another open-source web agent system; we configure it as a two-agent system: an AssistantAgent that interprets prompts and a MultimodalWebSurfer that executes browsing. It is faster and more controllable than Browser-Use, which makes it useful for both tasks. By default, AutoGen tends to summarize results directly without clicking through, so we enforce the same rule, visiting at least five pages.
@@
To collect metadata, we extend the web surfer into a custom LoggingWebSurfer that hooks Playwright APIs to record domains, timestamps, and IPs for all network events. Unlike GPT Researcher or Browser-Use, AutoGen logs not only primary domains but also auxiliary requests (ads, trackers, analytics), producing noisier but realistic traces that an actual network observer would see. We run AutoGen with GPT-4o for prompt recovery and Gemini 2.0 Flash for trait inference.
@@
OpenAI Deep Research API. As a proprietary reference point, we evaluate the OpenAI Deep Research API. Unlike the ChatGPT Deep Research interface, which engages in interactive follow-up questions, the API is stateless and relies entirely on instruction-rich inputs. To align its behavior with the UI, we provide -DR prompt variants (e.g., SESSION14-DR), rewritten to simulate the iterative refinement normally triggered by user interaction. Because its browsing occurs remotely, only provider connections are visible locally; we nonetheless analyze these traces to benchmark inference difficulty against a high-capability proprietary system.
@@
Rationale for Agent and Model Selection. We evaluate both research (GPT Researcher) and web agents (Browser-Use, AutoGen) to capture differences in browsing style and leakage. Each open-source agent is paired with both proprietary (GPT-4o, Gemini 2.0 Flash) and open-source (DeepSeek-V3, Mistral family) LLMs, reflecting realistic local deployments where users might expect stronger privacy protections. The OpenAI Deep Research API serves as a commercial baseline. Also, we deliberately exclude larger open-source models such as LLaMA 3.1 70B or DeepSeek-R1, as they require substantial computational resources and often bypass web search by relying on internal knowledge, making them less suitable for our local, browsing-driven evaluation. Together, these choices ensure our study captures attacker-visible metadata across diverse agents and for realistic local deployments, while also showing results against strong proprietary baselines.

### v3

@@ 7.1 Prompt Recovery
WRA vs. Human Attribution Analysis. To assess whether WRA traffic is distinguishable from human browsing, we clustered agent traces together with human sessions from SESSION14 using only timing, burstiness, and domain-transition features. WRAs consistently remain highly separable with 0.86–1.0 cluster purity across topics and 97.9% global classification accuracy. WRAs thus form a distinct, easily identifiable behavioral mode, exposing a richer and more attributable metadata footprint than humans.
ICL vs. Fine-Tuning Across Models. To motivate our choice of ICL over fine-tuning as the primary task-specific approach, we compare their performance across closed-source models (GPT-4o and GPT-4.1 nano) and open-source models (Llama 3.1 8B, Gemma 3 4B, and Qwen 2.5 32B) (Table 2). We create our final training set by combining the training samples from the prompt recovery datasets discussed in Section 6.1. For closed-source models, fine-tuning is performed through OpenAI’s fine-tuning API [openaiFinetuningAvailable] for 3 epochs with batch size 1. For open-source models, we apply LoRA [hu2022lora] with rank 16 and . Results show that ICL generally outperforms fine-tuning across nearly all metrics, though models like Llama occasionally favor fine-tuning for functional and domain equivalence. We attribute this to limited sample diversity, which likely causes overfitting in fine-tuned models.
@@ 7.2 Trait Inference
Figure 5 compares similarity scores for selected and unselected traits. Selected traits typically exceed 0.7, with some personas (012, 024, 037, 045) surpassing 0.8, indicating consistent high-risk leakage [cann2025using]. Unselected traits generally fall below but often remain above 0.5, showing moderate exposure even when traits are not explicitly embedded. In a few cases (007, 025, 041), unselected traits approach or exceed selected ones, confirming leakage extends beyond explicitly mentioned attributes.
@@ 8 Defense and Realistic Deployment
If a task requires visiting a unique domain, the defense should hide that interaction via VPNs or randomized background queries. By contrast, if the information can be obtained without distinctive traces, blocking is preferable; for example, by querying an LLM directly or gathering information from Wikipedia. We thus identify two conceptual mitigations for metadata-based inference attacks: (1) hiding traces, where sensitive activity must occur but is camouflaged with plausible noise; and (2) blocking traces, where the activity is avoided by relying on alternative, less revealing sources.
@@ 8.1 Hiding via Decoy Prompts
Inspired by TrackMeNot [nissenbaum2009trackmenot], for each real prompt the system generates and executes additional decoy prompts that remain within the same topical space but differ in framing, intent, or implied traits. Because OBELS evaluates whether reconstructions preserve intent, domain type, and entity, breaking this alignment is critical; otherwise, an adversary can still recover the user’s functional behavior despite wording changes. For instance, a prompt about “recognizing depression and seeking medical help for a friend” can be paired with a decoy like “compare the effectiveness of yoga versus meditation for improving sleep quality”, which shifts both domain and entity, lowering OBELS score and obscuring the true goal. This ambiguity in observed traffic undermines an adversary’s ability to infer the user’s true objective, attributes, or target entity. We explore a lightweight approach using LLM-based decoy generation and a trait-aware variant that leverages virtual personas to introduce long-term behavioral divergence.
@@ 8.2 Blocking via Alternative Sources
A complementary defense is to prevent sensitive traces from arising in the first place. When a user’s objective can be met without contacting uniquely identifying domains, the agent can instead draw from large multipurpose repositories (e.g., Wikipedia, StackExchange, Reddit) or from the LLM’s internal knowledge. By retrieving information from resources that serve diverse intents or by leveraging the model’s internal knowledge, the resulting traffic no longer maps cleanly to a specific user goal.
From the adversary’s perspective, this strategy effectively reduces visibility by collapsing domain diversity, yielding traces similar to those observed by weaker observers that see limited or non-discriminative domain information. This “blocking” approach therefore both mitigates leakage and illustrates how reduced observability weakens inference, reducing the adversary’s ability to associate domain visits with private attributes.
@@
Table 10 and Table 11 together show that visibility of browsing traces strongly shapes leakage but has little effect on utility. Utility is how much the WRA’s output helps a human reader reach their research goals. We measure it with an LLM using the prompt template. URL-level traces leak substantially more than domain-only traces, making prompt inference easier and more accurate. Reducing visibility (e.g., observing only 80% or 60% of domains) lowers leakage but still exposes useful information. Adding timing metadata to domains yields no consistent gains, suggesting that domain identity is more informative than temporal order. In contrast, the utility of the generated reports remains relatively stable across visibility levels down to 10%. However, beyond this threshold, utility begins to drop, revealing diminishing returns from additional domain coverage. This suggests that many extra domains contribute little to report quality while disproportionately increasing exposure to adversaries. Thus, WRA behaviors that aggressively visit many domains are especially problematic: they create substantial leakage risks while providing only marginal gains in utility.
@@ 9 Discussion and Conclusion
Privacy and Security Implications. Our results show that metadata leakage is an inherent risk for locally deployed web and research agents. We find that design choices—including domain exploration breadth, backbone model capacity, and orchestration logic—directly affect the amount of information exposed. Substituting locally deployable models does not eliminate leakage: domain-level traffic is structurally unavoidable. This shifts the attack vector from content to behavior. Whereas prompt injection and data exfiltration could be mitigated through isolation [tsai2025contextual], network-level traces cannot be trivially suppressed. At scale, such leakage enables profiling by advertisers, ISPs, or state actors, raising unique questions about how to balance agent utility with user privacy.
Defense Effectiveness and Limitations. Our study has several limitations. We rely on proxy datasets and synthetic personas with limited session counts; longer-term aggregation or multimodal attacks may amplify risks. Defense evaluation is confined to decoys and blocking, and cost/latency trade-offs are only partially measured. Nonetheless, our results show that decoy prompts and trait-conflicting personas reduce both prompt recovery and trait inference accuracy, while blocking strategies help when tasks can rely on broad, non-unique sources (e.g., Wikipedia) without sacrificing utility. Yet defenses remain partial: adversaries still recover useful signals, and decoys introduce overhead in latency, bandwidth, and behavioral realism. As with prior obfuscation systems, mitigation is meaningful but incomplete; full network-level protections (e.g., VPNs or anonymity systems for agents [pham2024proxygpt]) remain the only robust safeguard.
Future Directions and Open Challenges. Future work should explore hybrid defenses that combine obfuscation, blocking, and timing perturbations; study long-term aggregation attacks across extended user histories; and integrate formal privacy guarantees such as differential privacy or traffic-analysis resistance directly into agent design. Multi-modal agents introduce additional risks, for example, as passed multi-modal content could reveal even more information [schuster2017beauty]. Overall, progress requires viewing AI agents and the web as a coupled ecosystem rather than orthogonal components.
@@ Acknowledgments
The work was partially supported by Schmidt Sciences SAFE-AI program and by the NSF grants 2333965 and 2131910.
@@ Appendix A Ethical Considerations
Our research examines how WRAs can inadvertently leak sensitive information through metadata such as domain access patterns, timing, and interaction traces. While our goal is to strengthen privacy protections for agent-based systems, we recognize the potential for harmful use if such inference techniques were deployed maliciously.
Stakeholders. We identify four stakeholder groups affected by this research: (1) End-users of WRAs, who may have personal traits, goals, or intentions inferred without their awareness. This risk is not uniform; individuals in politically sensitive contexts, marginalized communities, or high-surveillance environments may face disproportionately higher harm. (2) Developers and framework maintainers, who may need to modify or harden system architectures, logging defaults, and data retention practices in response to identified risks. (3) Organizations deploying agents (e.g., educational platforms, customer service systems), which may unknowingly violate privacy expectations or regulatory obligations if leakage is not addressed. (4) The broader public, which has a collective interest in ensuring that emerging agent ecosystems evolve in privacy-preserving and autonomy-respecting directions.
Mitigating Potential Harm. To reduce risks to end-users, we designed our study so that no real user data were collected or analyzed; all experiments use public benchmark datasets (FEDWEB13, SESSION14, DD16) and synthetic personas, ensuring that no identifiable individuals can be profiled or re-identified. To support developers and framework maintainers, we present concrete mitigation strategies, such as trace obfuscation and domain suppression, that can be incorporated into WRA architectures to reduce metadata leakage. For organizations that deploy agents in educational, corporate, or customer-service environments, we do not release any tools for live traffic monitoring, and the accompanying code is limited to offline, dataset-based replay experiments, preventing misuse in operational settings. Finally, by identifying these risks early and providing actionable defenses, the work serves broader public interests by promoting privacy-preserving WRA design, rather than end-user surveillance or behavioral profiling as agents become widely deployed.
Researcher Well-being and Responsibilities. Our research did not involve direct interaction with online communities, covert participation, or exposure to disturbing or sensitive content. However, we continuously discussed the dual-use concerns throughout the project and made design choices aimed at minimizing the potential for misuse, for example, limiting our evaluation to public datasets and not designing tools that can be directly applied to live user traffic.
Societal Impact and Justification. As WRAs become integrated into productivity tools, classrooms, workplaces, and personal browsing workflows, metadata-based inference risks could scale rapidly and invisibly. By identifying and characterizing this attack vector under controlled conditions, we enable developers, regulators, and practitioners to recognize and address these vulnerabilities early, before they are exploited in real-world deployments. By focusing on public data, transparent methodology, and concrete mitigation guidance, we ensure that the research advances privacy-preserving agent design while minimizing the likelihood of harmful use.
@@ Appendix B Open Science
Implementations, evaluation code, prompts, and the synthetic datasets derived from our study will be made publicly available. Code and prompts available at: https://github.com/umass-aisec/wra.
@@ C.1 Dataset Details
Prompt Recovery. For all TREC datasets used in the prompt recovery task, we create -DR variants by inputting each original prompt with suggested_rewriting_prompt[alwell2025introduction], to GPT-4.1. The rewritten prompts are used for the Deep Research API runs to match its prompt handling, and all evaluations are scored against the original, unmodified prompts.
TREC FedWeb 2013 Topics contains 50 topics originally designed to evaluate federated web search. We concatenate the <description> and <narrative> fields to form a single prompt, and rewritten with GPT-4.1 (FEDWEB13-DR).
TREC Session 2014 Topics contains 60 realistic user topic descriptions spanning domains such as travel, health, and education. We use the original description fields as user prompts and their rewritten counterparts (SESSION14-DR).
TREC Dynamic Domain 2016 Truth Data provides 53 topics focused on interactive information retrieval in evolving domains such as medicine and cybersecurity. We concatenate the <description> and <narrative> fields and rewrite them with GPT-4.1 (DD16-DR).
@@
Trait Inference Personas. We use SynthLabsAI/PERSONA_subset to simulate benign users who use a web agent over time. The dataset includes 997 synthetic user profiles annotated with 32 traits in four categories (demographic, occupational, psychographic, and behavioral). We adopt a type-aware scoring for each trait for evaluation:
@@
Ordinal traits (e.g., big five scores) are mapped to integer levels and scored by normalized distance on a 5-point scale. Closer ratings yield higher similarity scores.
Categorical traits (e.g., gender, race) use exact match for single-token values; multi-word categories are scored using SBERT score to account for semantic equivalence.
@@
In our experiments, we sample 50 personas. For each, including those reserved for ICL examples, we randomly select 5 traits and use GPT-4o () to generate 21–35 trait-revealing prompts that explicitly or implicitly embed the traits. The produced traces serve as the input for trait inference.
| Task | ICL Shots | Output Format | Prompt Details |
| Prompt Recovery | 5 | Natural language query | Trace prompt |
| Trait Inference | 3 | Structured list | Trace trait list |
@@ C.2 Agent and LLM Backbone Choice
When browsing is executed locally, encrypted outbound traffic remains visible to an on-path adversary; when browsing is server-side (e.g., OpenAI’s Operator), only connections to the provider’s infrastructure are exposed. We collect traces using both open-source and proprietary WRAs.
@@
GPT Researcher [elovic2023gptresearcher]. GPT Researcher is a research agent designed for multi-hop investigation that decomposes prompts into sub-questions, gathers evidence, and synthesizes reports. We run it in report_type=‘‘deep’’ mode, enabling multi-turn exploration across sources. GPT Researcher employs different LLMs for sub-tasks: FAST_LLM for lightweight operations (e.g., GPT-4o-mini, DeepSeek-V3-0324), SMART_LLM for reasoning and report generation (e.g., GPT-4.1, Mistral-7B-Instruct v0.3), and STRATEGIC_LLM for higher-level planning (e.g., o4-mini, Mistral-Nemo-12B). While the default configuration uses GPT-4, we also evaluate local LLM combinations to assess leakage outside proprietary ecosystems.
@@
Browser-Use [browser_use2024]. Browser-Use is an open-source agent that automates browsing through a local browser. By default, it issues Google queries via search_web, which frequently triggers reCAPTCHA; we therefore configure it to use Bing with automatic fallback. We further modify its system prompt to require at least five distinct page visits before summarization, preventing it from shortcutting tasks with prior knowledge or superficial searches. After each run, we extract domains and timestamps from the browser’s recorded history to construct ordered traces. This setup makes Browser-Use suitable for evaluating how web agents’ lightweight browsing patterns leak information.
@@
AutoGen [wu2023autogen]. We configure AutoGen as a two-agent system consisting of an AssistantAgent and a MultimodalWebSurfer. Compared to Browser-Use, AutoGen is faster and more controllable, making it useful for both tasks. By default, AutoGen tends to summarize results directly without clicking through, so we enforce the same rule, visiting at least five pages. To collect metadata, we extend the web surfer into a custom LoggingWebSurfer that hooks Playwright APIs to record domains, timestamps, and IPs for all network events. Unlike GPT Researcher or Browser-Use, AutoGen logs not only primary domains but also auxiliary requests (ads, trackers, analytics), producing noisier but realistic traces that an actual network observer would see. We run AutoGen with GPT-4o for prompt recovery and Gemini 2.0 Flash for trait inference.
@@
OpenAI Deep Research API. As a proprietary baseline, we evaluate the OpenAI Deep Research API. Unlike the interactive ChatGPT Deep Research interface, the API is stateless and relies entirely on instruction-rich inputs. To align its behavior with the UI, we provide -DR prompt variants (e.g., SESSION14-DR), rewritten to simulate the iterative refinement that would have been normally triggered by user interaction. Because browsing occurs remotely, only provider-level connections are visible locally; we nonetheless analyze these traces to benchmark inference difficulty against a high-capability proprietary system.
@@
Rationale for Agent and Model Selection. We evaluate both research-oriented (GPT Researcher) and web-oriented (Browser-Use, AutoGen) agents to capture variation in browsing behavior and leakage. Each open-source agent is paired with both proprietary (GPT-4o, Gemini 2.0 Flash) and open-source (DeepSeek-V3, Mistral family) LLMs, reflecting realistic local deployments where users may expect stronger privacy guarantees. The Deep Research API serves as a commercial baseline. We deliberately exclude larger open-source models (e.g., LLaMA 3.1 70B, DeepSeek-R1), as they require substantial computational resources and often bypass web search by relying on internal knowledge, making them less suitable for our browsing-driven evaluation. Together, these choices ensure coverage of attacker-visible metadata across diverse agents and realistic local deployment settings.